Class Pulses Part 4 -Class.Pulses and Group Pulses e mean, median, mode, range, and standard deviation for the data. Us e the Use the data on the white sheet on the left to complete PART 6. Show formula used to create the classes and complete all of part 6
Answer #1

Class pulses : 75,77,69,72,74,56,59,69,71,72,77,68,54,89,84,80,74,65,68,73,75,93,79,85,71,65,78,80,56

Maximum = 93

Minimum = 54

Class width = Max - Min/ No. of classes = 93 -54/ 5 = 7.8

So first class= 54 - 61.8 and then so on

Class Interval Frequency

Relative Frequency

54-61.8 4 0.14
61.8-69.6 5 0.17
69.6-77.4 11 0.38
77.4-85.2 7 0.24
85.2-93 2 0.07
Total 29 1

Relative Frequency = Frequency / Total Frequency

3) . M is the midpoint of class interval and f is frequency

Classes M f M · f M2 · f
54-61.8 57.9 4 231.6 13409.64
61.8-69.6 65.7 5 328.5 21582.45
69.6-77.4 73.5 11 808.5 59424.75
77.4-85.2 81.3 7 569.1 46267.83
85.2-93 89.1 2 178.2 15877.62
Sum = 29 2115.9 156562.29

Now Frome the table

Mean = Sum of M*f / n = 2115.9/29 = 72.962

Variance = s2 = 1/N sum(m2f) - Mean2 = 1/29*156562.29 - 72.9622  = 75.24

Standard Deviation = s = (75.24)1/2 = 8.674

To find skewness where y is midpoint so m and n is no. of observation

skewness (n - 1)3

Classes M frequency (M-x)3*f
54-61.8 57.9 4 -13668.09287
61.8-69.6 65.7 5 -1914.867544
69.6-77.4 73.5 11 1.712929592
77.4-85.2 81.3 7 4057.735293
85.2-93 89.1 2 8405.80148
Sum = 29 -3117.710714

Since our sum is negative our data is left-skewed.